Determine if each function is a polynomial function or not. Determine the leading term, the constant term, and the degrees of each polynomial function.



  1. p(z) = z2 + 3z + 2017
  2. q(x) = 3x/2 + 4√x
  3. r(x) = x(x - 1) (x -2)2
  4. f(s) = 2017-3 + 2s-4 - 3s8
  5. A =  πr2
  6. A(e) = 6e2
1
Expert's answer
2021-01-29T12:52:05-0500

A polynomial in the variable "x" is an expression of the form


"a_nx^n+a_{n-1}x^{n-1}+...+a_1x+a_0"


where "a_0,a_1,...a_n" are real numbers, and "n" is a nonnegative integer. 


1. "p(z)=z^2+3z+2017"

"p(z)" is a polynomial function in the variable "z"

Leading term is "z^2," constant term is "2017," degree is "2."


2. "q(x)=3x/2+\sqrt[4]{x}"

"q(x)" is not a polynomial function in the variable "x"


3. "r(x)=x(x-1)(x-2)^2"

"r(x)" is a polynomial function in the variable "x"

Leading term is "x^4," constant term is "0," degree is "4."


4. "f(s)=2017^{-3}+2s^{-4}-3s^8"

"f(s)" is not a polynomial function in the variable "s"


5. "A=\pi r^2"

"A(r)" is a polynomial function in the variable "r"

Leading term is "\pi r^2," constant term is "0," degree is "2."


6. "A(e)=6e^2"

"A(e)" is a polynomial function in the variable "e"

Leading term is "6e^2," constant term is "0," degree is "0."


If "e" is a mathemical constant (Euler's number), then A is number and is a polynomial function of degree 0 in the any variable .

Leading term is "6e^2," constant term is "6e^2," degree is "0."